# Hierarchical Time Series Forecasting with Robust Reconciliation

### v1

@@ Appendix A Full results of numerical experiments
total $1726.07$ $3680.00$ $1726.07$ $1906.48$ $1204.56$ $\underline{1101.69}$
NSW $814.62$ $1153.92$ $\underline{390.45}$ $709.89$ $556.17$ $765.66$
NT $135.97$ $136.13$ $177.71$ $62.33$ $101.17$ $\underline{61.77}$
QLD $633.93$ $691.06$ $581.55$ $503.91$ $264.39$ $\underline{262.21}$
SA $226.56$ $291.26$ $105.01$ $\underline{98.52}$ $99.75$ $100.02$
TAS $148.29$ $148.68$ $185.51$ $\underline{86.92}$ $96.55$ $95.52$
VIC $642.69$ $938.29$ $697.81$ $553.19$ $\underline{475.28}$ $775.65$
WA $362.15$ $481.64$ $513.88$ $272.77$ $\underline{233.58}$ $514.91$
NSW_ACT $146.72$ $146.72$ $86.05$ $\underline{79.23}$ $83.41$ $88.06$
NSW_Nth $158.32$ $158.32$ $110.01$ $93.05$ $99.64$ $\underline{67.53}$
NSW_Sth $145.49$ $145.49$ $128.61$ $\underline{90.52}$ $104.94$ $186.33$
NSW_Metro $370.82$ $370.82$ $\underline{192.91}$ $300.52$ $178.58$ $329.89$
NSW_Nth_Coast $265.32$ $265.32$ $145.86$ $193.64$ $\underline{162.47}$ $211.55$
NSW_Sth_Coast $131.95$ $131.95$ $165.04$ $64.79$ $\underline{54.05}$ $130.65$
NT_Central $87.08$ $87.08$ $102.50$ $\underline{48.88}$ $69.71$ $56.96$
NT_Nth_Coast $58.57$ $58.57$ $81.15$ $\underline{36.06}$ $43.98$ $73.95$
QLD_Metro $432.75$ $432.75$ $278.40$ $389.27$ $243.99$ $\underline{211.76}$
QLD_Central_Coast $62.53$ $62.53$ $78.83$ $\underline{42.89}$ $48.47$ $45.56$
QLD_Inland $174.46$ $174.46$ $194.94$ $131.45$ $\underline{78.24}$ $98.68$
QLD_Nth_Coast $101.77$ $101.77$ $183.26$ $73.57$ $\underline{70.65}$ $127.87$
SA_Metro $134.08$ $134.08$ $41.14$ $84.94$ $45.04$ $\underline{40.84}$
SA_Inland $87.19$ $87.19$ $\underline{36.99}$ $43.59$ $42.36$ $38.44$
SA_West_Coast $31.63$ $31.63$ $33.36$ $35.80$ $\underline{21.76}$ $25.81$
SA_Sth_Coast $66.34$ $66.34$ $82.00$ $\underline{42.47}$ $46.66$ $92.30$
TAS_Nth_East $58.96$ $58.96$ $70.44$ $\underline{35.35}$ $39.60$ $43.19$
TAS_Sth $61.49$ $61.49$ $84.42$ $\underline{47.51}$ $48.14$ $53.70$
TAS_Nth_West $44.88$ $44.88$ $42.50$ $\underline{30.26}$ $33.01$ $66.44$
VIC_Nth_West $124.89$ $124.89$ $86.11$ $76.51$ $\underline{74.90}$ $87.18$
VIC_Nth_East $232.47$ $232.47$ $\underline{108.39}$ $153.51$ $116.29$ $132.61$
VIC_Metro $385.23$ $385.23$ $397.61$ $307.86$ $\underline{238.76}$ $386.30$
VIC_East_Coast $142.78$ $142.78$ $164.21$ $\underline{87.36}$ $97.22$ $180.77$
VIC_West_Coast $101.05$ $101.05$ $122.62$ $\underline{43.97}$ $50.62$ $56.05$
WA_West_Coast $275.32$ $275.32$ $364.37$ $\underline{214.80}$ $225.39$ $419.57$
WA_Sth $101.13$ $101.13$ $63.93$ $\underline{41.77}$ $42.26$ $70.69$
WA_Nth $135.48$ $135.48$ $134.23$ $63.10$ $\underline{48.77}$ $54.08$
@@
total $2341.30$ $2672.04$ $2341.30$ $2369.42$ $2257.15$ $\underline{2069.26}$
CA $1235.18$ $1390.02$ $\underline{473.17}$ $1243.65$ $1192.38$ $1115.32$
TX $588.46$ $588.94$ $1363.35$ $567.46$ $525.10$ $\underline{496.01}$
WI $551.08$ $700.38$ $594.97$ $567.29$ $551.10$ $\underline{466.53}$
CA_1 $218.13$ $218.13$ $509.39$ $\underline{183.49}$ $196.72$ $188.12$
CA_2 $702.68$ $702.68$ $626.25$ $666.09$ $555.67$ $\underline{493.65}$
CA_3 $202.95$ $202.95$ $371.52$ $\underline{166.36}$ $192.27$ $192.00$
CA_4 $266.87$ $266.87$ $\underline{153.50}$ $230.28$ $248.95$ $241.55$
TX_1 $144.27$ $144.27$ $495.96$ $137.11$ $122.65$ $\underline{110.52}$
TX_2 $205.94$ $205.94$ $504.96$ $198.77$ $187.25$ $\underline{180.26}$
TX_3 $239.98$ $239.98$ $362.43$ $232.81$ $215.85$ $\underline{206.92}$
WI_1 $270.88$ $270.88$ $\underline{59.44}$ $226.51$ $192.03$ $142.00$
WI_2 $181.19$ $181.19$ $171.11$ $137.77$ $129.79$ $\underline{98.80}$
WI_3 $251.52$ $251.52$ $456.19$ $\underline{206.62}$ $233.63$ $231.29$
@@
total $2476.17$ $2813.26$ $2476.17$ $2505.82$ $2390.48$ $\underline{2195.19}$
CA $1312.65$ $1459.66$ $\underline{500.67}$ $1318.29$ $1264.27$ $1190.99$
TX $616.51$ $617.08$ $1385.58$ $594.22$ $549.17$ $\underline{517.10}$
WI $585.71$ $743.75$ $634.64$ $602.89$ $586.51$ $\underline{498.61}$
CA_1 $229.50$ $229.50$ $516.40$ $\underline{193.23}$ $206.75$ $197.65$
CA_2 $756.67$ $756.67$ $631.13$ $723.15$ $621.26$ $\underline{573.69}$
CA_3 $209.95$ $209.95$ $376.76$ $\underline{172.81}$ $198.88$ $197.87$
CA_4 $277.57$ $277.57$ $\underline{163.07}$ $240.85$ $259.08$ $250.52$
TX_1 $152.46$ $152.46$ $503.13$ $144.98$ $129.80$ $\underline{116.76}$
TX_2 $214.92$ $214.92$ $511.64$ $207.35$ $195.32$ $\underline{187.26}$
TX_3 $251.75$ $251.75$ $371.40$ $244.08$ $226.39$ $\underline{215.27}$
WI_1 $289.03$ $289.03$ $\underline{90.64}$ $241.61$ $205.28$ $152.04$
WI_2 $194.00$ $194.00$ $183.29$ $149.06$ $141.13$ $\underline{112.22}$
WI_3 $264.39$ $264.39$ $464.58$ $\underline{217.38}$ $245.68$ $242.46$
@@
grid $39.615$ $41.587$ $39.615$ $40.410$ $44.313$ $\underline{35.362}$
S1 $18.900$ $19.164$ $24.268$ $\underline{18.844}$ $19.059$ $18.992$
S2 $32.251$ $35.129$ $22.439$ $31.545$ $32.255$ $\underline{20.321}$
S11 $7.743$ $8.514$ $17.003$ $7.852$ $7.641$ $\underline{7.629}$
S12 $11.359$ $\underline{11.358}$ $14.274$ $11.361$ $11.499$ $11.363$
S21 $25.096$ $25.621$ $22.450$ $24.633$ $22.938$ $\underline{13.200}$
S22 $8.353$ $9.748$ $10.231$ $8.242$ $9.617$ $\underline{8.137}$
S11_1 $1.093$ $1.093$ $1.867$ $0.789$ $\underline{0.485}$ $0.491$
S11_2 $2.261$ $2.261$ $4.386$ $2.620$ $2.610$ $\underline{2.024}$
S11_3 $\underline{1.756}$ $\underline{1.756}$ $2.318$ $2.123$ $2.183$ $1.930$
S11_4 $3.007$ $3.007$ $3.340$ $2.966$ $\underline{2.908}$ $2.991$
S11_5 $5.959$ $5.959$ $15.835$ $5.688$ $5.730$ $\underline{5.201}$
S11_6 $2.624$ $2.624$ $3.617$ $2.733$ $2.774$ $\underline{2.617}$
S12_1 $\underline{4.556}$ $\underline{4.556}$ $8.824$ $4.590$ $4.742$ $5.093$
S12_2 $1.776$ $1.776$ $\underline{1.667}$ $1.738$ $2.038$ $2.181$
S12_3 $4.502$ $4.502$ $7.805$ $\underline{4.454}$ $5.275$ $4.640$
S12_4 $4.230$ $4.230$ $6.363$ $4.264$ $4.272$ $\underline{3.132}$
S12_5 $4.332$ $4.332$ $4.246$ $4.370$ $\underline{3.950}$ $5.448$
S12_6 $2.504$ $2.504$ $\underline{2.156}$ $2.461$ $2.889$ $2.669$
S21_1 $1.339$ $1.339$ $2.927$ $1.442$ $\underline{1.172}$ $1.438$
S21_2 $1.121$ $1.121$ $1.132$ $1.173$ $\underline{1.109}$ $1.265$
S21_3 $0.876$ $0.876$ $2.786$ $0.849$ $\underline{0.844}$ $1.361$
S21_4 $2.748$ $2.748$ $10.603$ $2.810$ $\underline{2.712}$ $3.279$
S21_5 $6.577$ $6.577$ $\underline{4.226}$ $6.413$ $7.174$ $5.530$
S21_6 $17.532$ $17.532$ $18.858$ $17.367$ $13.743$ $\underline{9.232}$
S22_1 $2.160$ $2.160$ $5.757$ $\underline{1.823}$ $2.475$ $2.315$
S22_2 $1.965$ $1.965$ $7.783$ $\underline{1.696}$ $2.161$ $2.122$
S22_3 $3.879$ $3.879$ $9.010$ $\underline{3.688}$ $3.720$ $4.363$
S22_4 $1.099$ $1.099$ $1.980$ $1.207$ $\underline{1.097}$ $1.101$
S22_5 $1.220$ $1.220$ $1.179$ $1.079$ $1.262$ $\underline{1.072}$
S22_6 $1.190$ $1.190$ $2.002$ $\underline{0.995}$ $1.612$ $2.043$
@@
grid $64.441$ $66.057$ $64.441$ $65.156$ $67.813$ $\underline{61.076}$
S1 $30.936$ $\underline{30.400}$ $34.750$ $30.722$ $31.714$ $31.483$
S2 $39.334$ $41.592$ $32.570$ $38.789$ $39.311$ $\underline{30.959}$
S11 $12.054$ $\underline{11.937}$ $20.505$ $11.968$ $12.354$ $12.442$
S12 $19.044$ $19.041$ $\underline{17.878}$ $18.984$ $19.535$ $19.208$
S21 $28.566$ $28.995$ $27.339$ $28.181$ $26.775$ $\underline{19.874}$
S22 $11.779$ $13.269$ $12.713$ $11.676$ $13.129$ $\underline{11.617}$
S11_1 $1.172$ $1.172$ $2.092$ $0.883$ $\underline{0.655}$ $0.657$
S11_2 $2.795$ $2.795$ $5.071$ $3.111$ $3.103$ $\underline{2.552}$
S11_3 $\underline{2.009}$ $\underline{2.009}$ $2.772$ $2.347$ $2.403$ $2.160$
S11_4 $3.464$ $3.464$ $4.340$ $\underline{3.408}$ $3.607$ $3.858$
S11_5 $6.308$ $6.308$ $16.369$ $6.017$ $6.062$ $\underline{5.514}$
S11_6 $4.689$ $4.689$ $5.593$ $4.832$ $4.864$ $\underline{4.669}$
S12_1 $6.568$ $6.568$ $12.232$ $6.569$ $\underline{6.564}$ $6.622$
S12_2 $2.366$ $2.366$ $\underline{2.104}$ $2.332$ $2.599$ $2.717$
S12_3 $5.022$ $5.022$ $8.321$ $\underline{4.978}$ $5.736$ $5.140$
S12_4 $4.704$ $4.704$ $7.391$ $4.728$ $4.733$ $\underline{4.185}$
S12_5 $5.151$ $5.151$ $5.078$ $5.174$ $\underline{4.954}$ $5.972$
S12_6 $2.743$ $2.743$ $2.983$ $\underline{2.699}$ $3.104$ $2.899$
S21_1 $1.726$ $1.726$ $3.776$ $1.785$ $\underline{1.654}$ $1.782$
S21_2 $\underline{1.564}$ $\underline{1.564}$ $1.613$ $1.590$ $1.566$ $1.631$
S21_3 $1.028$ $1.028$ $3.708$ $0.998$ $\underline{0.992}$ $1.602$
S21_4 $4.410$ $4.410$ $11.256$ $4.364$ $4.455$ $\underline{4.228}$
S21_5 $7.195$ $7.195$ $\underline{5.089}$ $7.047$ $7.737$ $6.277$
S21_6 $18.540$ $18.540$ $19.633$ $18.392$ $15.176$ $\underline{11.590}$
S22_1 $2.547$ $2.547$ $6.196$ $\underline{2.209}$ $2.841$ $2.690$
S22_2 $2.387$ $2.387$ $8.250$ $\underline{2.055}$ $2.640$ $2.593$
S22_3 $6.343$ $6.343$ $10.262$ $6.134$ $5.684$ $\underline{5.568}$
S22_4 $1.324$ $1.324$ $2.462$ $1.454$ $\underline{1.308}$ $1.318$
S22_5 $1.713$ $1.713$ $1.498$ $1.500$ $1.753$ $\underline{1.495}$
S22_6 $1.455$ $1.455$ $2.422$ $\underline{1.226}$ $1.869$ $2.287$

### v2

@@ Appendix A Full results of numerical experiments
total $1726.07$ $3680.00$ $1726.07$ $1906.48$ $1204.56$ $\underline{895.21}$
NSW $814.62$ $1153.92$ $\underline{390.45}$ $709.89$ $556.17$ $767.23$
NT $135.97$ $136.13$ $177.71$ $\underline{62.33}$ $101.17$ $90.27$
QLD $633.93$ $691.06$ $581.55$ $503.91$ $\underline{264.39}$ $267.44$
SA $226.56$ $291.26$ $105.01$ $98.52$ $99.75$ $\underline{94.40}$
TAS $148.29$ $148.68$ $185.51$ $86.92$ $96.55$ $\underline{78.14}$
VIC $642.69$ $938.29$ $697.81$ $553.19$ $\underline{475.28}$ $725.78$
WA $362.15$ $481.64$ $513.88$ $272.77$ $\underline{233.58}$ $567.47$
NSW_ACT $146.72$ $146.72$ $86.05$ $\underline{79.23}$ $83.41$ $117.14$
NSW_Nth $158.32$ $158.32$ $110.01$ $93.05$ $99.64$ $\underline{73.21}$
NSW_Sth $145.49$ $145.49$ $128.61$ $\underline{90.52}$ $104.94$ $186.20$
NSW_Metro $370.82$ $370.82$ $192.91$ $300.52$ $\underline{178.58}$ $365.35$
NSW_Nth_Coast $265.32$ $265.32$ $\underline{145.86}$ $193.64$ $162.47$ $222.41$
NSW_Sth_Coast $131.95$ $131.95$ $165.04$ $64.79$ $\underline{54.05}$ $122.21$
NT_Central $87.08$ $87.08$ $102.50$ $48.88$ $69.71$ $\underline{41.73}$
NT_Nth_Coast $58.57$ $58.57$ $81.15$ $\underline{36.06}$ $43.98$ $81.02$
QLD_Metro $432.75$ $432.75$ $278.40$ $389.27$ $243.99$ $\underline{207.95}$
QLD_Central_Coast $62.53$ $62.53$ $78.83$ $\underline{42.89}$ $48.47$ $64.68$
QLD_Inland $174.46$ $174.46$ $194.94$ $131.45$ $\underline{78.24}$ $109.60$
QLD_Nth_Coast $101.77$ $101.77$ $183.26$ $73.57$ $\underline{70.65}$ $145.45$
SA_Metro $134.08$ $134.08$ $\underline{41.14}$ $84.94$ $45.04$ $43.86$
SA_Inland $87.19$ $87.19$ $\underline{36.99}$ $43.59$ $42.36$ $46.13$
SA_West_Coast $31.63$ $31.63$ $33.36$ $35.80$ $\underline{21.76}$ $31.26$
SA_Sth_Coast $66.34$ $66.34$ $82.00$ $\underline{42.47}$ $46.66$ $69.52$
TAS_Nth_East $58.96$ $58.96$ $70.44$ $\underline{35.35}$ $39.60$ $47.58$
TAS_Sth $61.49$ $61.49$ $84.42$ $\underline{47.51}$ $48.14$ $55.55$
TAS_Nth_West $44.88$ $44.88$ $42.50$ $30.26$ $33.01$ $\underline{25.17}$
VIC_Nth_West $124.89$ $124.89$ $86.11$ $76.51$ $\underline{74.90}$ $81.43$
VIC_Nth_East $232.47$ $232.47$ $108.39$ $153.51$ $116.29$ $\underline{106.73}$
VIC_Metro $385.23$ $385.23$ $397.61$ $307.86$ $\underline{238.76}$ $370.23$
VIC_East_Coast $142.78$ $142.78$ $164.21$ $\underline{87.36}$ $97.22$ $177.92$
VIC_West_Coast $101.05$ $101.05$ $122.62$ $\underline{43.97}$ $50.62$ $68.44$
WA_West_Coast $275.32$ $275.32$ $364.37$ $\underline{214.80}$ $225.39$ $446.61$
WA_Sth $101.13$ $101.13$ $63.93$ $\underline{41.77}$ $42.26$ $79.92$
WA_Nth $135.48$ $135.48$ $134.23$ $63.10$ $\underline{48.77}$ $66.16$
@@
total $2343.98$ $2659.41$ $2343.98$ $2370.30$ $2264.97$ $\underline{1898.06}$
CA $1237.39$ $1397.73$ $\underline{473.01}$ $1248.40$ $1195.54$ $993.95$
TX $586.12$ $587.64$ $1364.16$ $566.09$ $529.02$ $\underline{447.99}$
WI $551.73$ $682.22$ $595.33$ $563.94$ $551.06$ $\underline{460.93}$
CA_1 $217.28$ $217.28$ $509.66$ $182.14$ $196.40$ $\underline{169.73}$
CA_2 $707.61$ $707.61$ $626.00$ $670.27$ $556.24$ $\underline{426.06}$
CA_3 $205.31$ $205.31$ $371.79$ $\underline{167.98}$ $193.55$ $174.84$
CA_4 $268.06$ $268.06$ $\underline{153.50}$ $230.73$ $250.44$ $224.32$
TX_1 $139.57$ $139.57$ $496.24$ $132.39$ $122.58$ $\underline{95.30}$
TX_2 $210.50$ $210.50$ $505.23$ $203.52$ $190.62$ $\underline{161.45}$
TX_3 $239.12$ $239.12$ $362.70$ $232.15$ $216.86$ $\underline{193.64}$
WI_1 $266.21$ $266.21$ $\underline{60.84}$ $226.78$ $197.83$ $129.05$
WI_2 $178.76$ $178.76$ $171.12$ $139.94$ $\underline{135.31}$ $138.43$
WI_3 $238.70$ $238.70$ $456.46$ $199.22$ $220.13$ $\underline{195.16}$
@@
total $2479.68$ $2799.50$ $2479.68$ $2507.51$ $2398.58$ $\underline{2026.11}$
CA $1315.76$ $1466.86$ $\underline{502.11}$ $1323.84$ $1267.95$ $1076.92$
TX $613.93$ $615.60$ $1386.53$ $592.66$ $553.26$ $\underline{469.26}$
WI $586.42$ $724.49$ $635.82$ $599.93$ $586.67$ $\underline{491.14}$
CA_1 $228.58$ $228.58$ $516.72$ $191.94$ $206.43$ $\underline{178.76}$
CA_2 $761.58$ $761.58$ $630.90$ $727.71$ $623.03$ $\underline{518.46}$
CA_3 $212.33$ $212.33$ $377.04$ $\underline{174.67}$ $200.11$ $180.73$
CA_4 $278.59$ $278.59$ $\underline{163.45}$ $241.39$ $260.40$ $233.43$
TX_1 $147.41$ $147.41$ $503.43$ $139.73$ $129.54$ $\underline{101.01}$
TX_2 $219.32$ $219.32$ $511.96$ $211.67$ $198.63$ $\underline{168.66}$
TX_3 $250.99$ $250.99$ $371.73$ $243.43$ $227.41$ $\underline{202.47}$
WI_1 $283.85$ $283.85$ $\underline{90.82}$ $241.77$ $211.29$ $136.84$
WI_2 $191.39$ $191.39$ $183.70$ $151.72$ $\underline{147.07}$ $151.58$
WI_3 $252.62$ $252.62$ $464.93$ $211.08$ $233.04$ $\underline{207.25}$
@@
grid $39.471$ $41.091$ $39.471$ $40.343$ $43.814$ $\underline{35.068}$
S1 $18.902$ $19.298$ $24.210$ $\underline{18.845}$ $19.039$ $18.977$
S2 $32.668$ $34.938$ $22.365$ $31.639$ $32.098$ $\underline{19.997}$
S11 $7.783$ $8.606$ $16.944$ $7.880$ $7.647$ $\underline{7.629}$
S12 $11.358$ $11.356$ $14.279$ $11.362$ $11.437$ $\underline{11.348}$
S21 $24.839$ $25.529$ $22.372$ $24.545$ $22.737$ $\underline{12.900}$
S22 $8.388$ $9.680$ $10.263$ $8.347$ $9.649$ $\underline{8.115}$
S11_1 $1.087$ $1.087$ $1.864$ $0.783$ $\underline{0.478}$ $0.489$
S11_2 $2.265$ $2.265$ $4.394$ $2.627$ $2.666$ $\underline{2.000}$
S11_3 $\underline{1.710}$ $\underline{1.710}$ $2.321$ $2.078$ $2.142$ $1.877$
S11_4 $3.008$ $3.008$ $3.335$ $2.968$ $\underline{2.909}$ $2.994$
S11_5 $6.053$ $6.053$ $15.819$ $5.771$ $5.806$ $\underline{5.135}$
S11_6 $2.619$ $2.619$ $3.602$ $2.724$ $2.753$ $\underline{2.616}$
S12_1 $\underline{4.558}$ $\underline{4.558}$ $8.822$ $4.563$ $4.760$ $5.088$
S12_2 $1.758$ $1.758$ $\underline{1.667}$ $1.756$ $2.023$ $2.144$
S12_3 $4.519$ $4.519$ $7.814$ $\underline{4.515}$ $5.325$ $4.572$
S12_4 $4.263$ $4.263$ $6.355$ $4.268$ $4.233$ $\underline{3.121}$
S12_5 $4.499$ $4.499$ $4.249$ $4.504$ $\underline{4.169}$ $5.501$
S12_6 $2.505$ $2.505$ $\underline{2.161}$ $2.502$ $2.884$ $2.610$
S21_1 $1.359$ $1.359$ $2.935$ $1.470$ $\underline{1.174}$ $1.490$
S21_2 $1.124$ $1.124$ $1.135$ $1.180$ $\underline{1.106}$ $1.297$
S21_3 $0.872$ $0.872$ $2.793$ $0.848$ $\underline{0.841}$ $1.385$
S21_4 $2.751$ $2.751$ $10.583$ $2.812$ $\underline{2.745}$ $3.352$
S21_5 $6.562$ $6.562$ $\underline{4.230}$ $6.398$ $7.161$ $5.460$
S21_6 $17.541$ $17.541$ $18.843$ $17.377$ $13.768$ $\underline{9.224}$
S22_1 $2.173$ $2.173$ $5.764$ $\underline{1.866}$ $2.493$ $2.296$
S22_2 $1.928$ $1.928$ $7.793$ $\underline{1.687}$ $2.118$ $2.091$
S22_3 $3.880$ $3.880$ $8.993$ $3.711$ $\underline{3.702}$ $4.317$
S22_4 $1.100$ $1.100$ $1.983$ $1.198$ $\underline{1.093}$ $1.101$
S22_5 $1.225$ $1.225$ $1.180$ $1.088$ $1.266$ $\underline{1.066}$
S22_6 $1.168$ $1.168$ $1.998$ $\underline{0.994}$ $1.581$ $2.003$
@@
grid $64.301$ $65.685$ $64.301$ $65.102$ $67.498$ $\underline{60.821}$
S1 $30.940$ $\underline{30.341}$ $34.684$ $30.683$ $31.597$ $31.432$
S2 $39.665$ $41.439$ $32.503$ $38.863$ $39.198$ $\underline{30.729}$
S11 $12.023$ $11.953$ $20.456$ $\underline{11.951}$ $12.321$ $12.442$
S12 $19.037$ $18.966$ $\underline{17.871}$ $18.976$ $19.450$ $19.160$
S21 $28.347$ $28.919$ $27.276$ $28.105$ $26.616$ $\underline{19.665}$
S22 $11.817$ $13.200$ $12.733$ $11.781$ $13.164$ $\underline{11.596}$
S11_1 $1.166$ $1.166$ $2.088$ $0.877$ $\underline{0.649}$ $0.656$
S11_2 $2.800$ $2.800$ $5.079$ $3.117$ $3.152$ $\underline{2.526}$
S11_3 $\underline{1.968}$ $\underline{1.968}$ $2.775$ $2.304$ $2.365$ $2.111$
S11_4 $3.466$ $3.466$ $4.334$ $\underline{3.409}$ $3.596$ $3.862$
S11_5 $6.399$ $6.399$ $16.354$ $6.105$ $6.142$ $\underline{5.449}$
S11_6 $4.682$ $4.682$ $5.583$ $4.823$ $4.850$ $\underline{4.666}$
S12_1 $6.568$ $6.568$ $12.222$ $6.571$ $\underline{6.566}$ $6.621$
S12_2 $2.350$ $2.350$ $\underline{2.105}$ $2.348$ $2.586$ $2.688$
S12_3 $5.037$ $5.037$ $8.329$ $\underline{5.034}$ $5.782$ $5.077$
S12_4 $4.726$ $4.726$ $7.384$ $4.730$ $4.707$ $\underline{4.183}$
S12_5 $5.247$ $5.247$ $5.082$ $5.252$ $\underline{5.062}$ $6.016$
S12_6 $2.744$ $2.744$ $2.989$ $\underline{2.741}$ $3.099$ $2.844$
S21_1 $1.739$ $1.739$ $3.783$ $1.800$ $\underline{1.655}$ $1.811$
S21_2 $1.566$ $1.566$ $1.614$ $1.592$ $\underline{1.560}$ $1.648$
S21_3 $1.022$ $1.022$ $3.715$ $0.996$ $\underline{0.993}$ $1.625$
S21_4 $4.406$ $4.406$ $11.238$ $4.360$ $4.415$ $\underline{4.228}$
S21_5 $7.181$ $7.181$ $\underline{5.095}$ $7.034$ $7.725$ $6.216$
S21_6 $18.548$ $18.548$ $19.619$ $18.400$ $15.199$ $\underline{11.585}$
S22_1 $2.559$ $2.559$ $6.203$ $\underline{2.256}$ $2.858$ $2.672$
S22_2 $2.339$ $2.339$ $8.260$ $\underline{2.046}$ $2.590$ $2.557$
S22_3 $6.345$ $6.345$ $10.247$ $6.159$ $5.731$ $\underline{5.548}$
S22_4 $1.326$ $1.326$ $2.465$ $1.443$ $\underline{1.310}$ $1.320$
S22_5 $1.717$ $1.717$ $1.497$ $1.522$ $1.758$ $\underline{1.484}$
S22_6 $1.424$ $1.424$ $2.417$ $\underline{1.226}$ $1.838$ $2.248$